OpenStudy (anonymous):

you have $26, and your brother has $38. You are saving $10 per day and your brother is saving $6 per day. After how many days will you and your brother have the same amount of money? The y-intercept is the amount you start with (use y = mx + b)

OpenStudy (jtug6):

ill help after i finish this game, few mins :p

OpenStudy (jtug6):

ok first things first youre gonna wanna write it out in it's form y = mx + b for whats given so: m = slope or rate and b = what youre starting with so y1 = 10x + 26 and y2 = 6x + 38 and we want to know when x is equal to each other so we say 10x + 26 = 6x + 38 can you solve for x?

OpenStudy (anonymous):

x = 3

OpenStudy (jtug6):

Yep. So now you know that x = 3 or 3 days later you have the same amount of $

OpenStudy (anonymous):

Oh ok thank you so much! :) That actually makes a lot of sense.

OpenStudy (anonymous):

@jtug6 can you help with one more question?

OpenStudy (jtug6):

umm sure

OpenStudy (anonymous):

A student scored 73 and 89 on his first two quizzes. Use a compound inequality to find the possible values for a third quiz score that would give him an average between 80 and 85, inclusive. Hint: Inclusive means >_ and <_. Make sure to find the average of the three test scores (use x to represent the third test score).

OpenStudy (anonymous):

@jtug6

OpenStudy (jtug6):

so basically it's just 80 <= (73+89+x)/3 <= 85 because to find avg you need to add all 3 and divide by 3

OpenStudy (jtug6):

or <_ >_ instead of = :P

OpenStudy (jtug6):

Was that it?